### v1

@@ 3 General Inference Framework
We now introduce the general inference framework that we are interested in, nesting the Manski bound example of the previous section as a special case. After introducing the general framework, we describe several additional example applications that fall within this framework.
We are interested in performing inference on a parameter $W(d)$ that is indexed by a finite set $d\in\mathcal{D}\equiv\{0,\ldots,T\}$ for some $T>0$. The index $d$ may correspond to a particular treatment, treatment allocation rule or policy, depending upon the application. We assume that $W(d)$ belongs to an identified set taking a particular interval form that is common to many applications of interest.
@@
For all $d\in\{0,\ldots,T\}$ and an unknown finite-dimensional parameter $p$,
@@
    $L(d)\equiv\max_{j\in\{1,\ldots,J_{L}\}}\{\tilde{\ell}_{d,j}+\ell_{d,j}p\}\leq W(d)$ for some fixed and known $J_{L}$, ~$\tilde{\ell}_{d,1},\ldots,\tilde{\ell}_{d,J_{L}}$ and nonzero $\ell_{d,1},\ldots,\ell_{d,J_{L}}$ such that $\ell_{d,j}\neq\ell_{d,j^{\prime}}$ for $j\neq j^{\prime}$.
@@
    $U(d)\equiv\min_{j\in\{1,\ldots,J_{U}\}}\{\tilde{u}_{d,j}+u_{d,j}p\}\geq W(d)$ for some fixed and known $J_{U}$, $\tilde{u}_{d,1},\ldots,\tilde{u}_{d,J_{U}}$ and nonzero $u_{d,1},\ldots,u_{d,J_{U}}$ such that $u_{d,j}\neq u_{d,j^{\prime}}$ for $j\neq j^{\prime}$.
The lower and upper endpoints of identified sets for the welfare, average potential outcome or ATE of a treatment $d$ typically take the form of $L(d)$ and $U(d)$ in cases for which (sequences of) outcomes, treatments and instruments are discrete.
In the setting of this paper, a researcher’s interest in $W(d)$ arises when $d$ belongs to a set $\widehat{\mathcal{D}}\subset\{0,\ldots,T\}$ that is estimated from a sample of $n$ observations. It is often the case that $\widehat{\mathcal{D}}$ is an estimate of $\mathcal{D}^{*}$. This set could correspond to an estimated set of optimal treatments or policies or other data-dependent index sets of interest. The estimated set is determined by an estimator $\hat{p}$ of the finite-dimensional parameter $p$ that determines the bounds on $W(d)$ according to Assumption 3.1. Let
$\hat{j}_{L}(d)=\operatorname*{argmax\;}_{j\in\{1,\ldots,J_{L}\}}\{\tilde{\ell}_{d,j}+\ell_{d,j}\hat{p}\},\quad\hat{j}_{U}(d)=\operatorname{argmin}_{j\in\{1,\ldots,J_{U}\}}\{\tilde{u}_{d,j}+u_{d,j}\hat{p}\}$
so that the estimated lower and upper bounds for $W(d)$ are equal to ~$\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p}$ and $\tilde{u}_{d,\hat{j}_{U}(d)}+u_{d,\hat{j}_{U}(d)}\hat{p}$. We work under the high-level assumption that the event that (i) a parameter is in the set of interest, (ii) the estimated bounds on $W(d)$ are realized at a given value and (iii) an additional random vector is realized at any given value can be written as a polyhedron in $\hat{p}$.
@@
    For some finite-valued random vector $\hat{\gamma}_{L}(d)$, some fixed and known matrix $A^{L}(d,j_{L}^{*},\gamma_{L}^{*})$ and some fixed and known vector $c^{L}(d,j_{L}^{*},\gamma_{L}^{*})$, $d\in\widehat{\mathcal{D}}$, $\hat{j}_{L}(d)=j_{L}^{*}$ and $\hat{\gamma}_{L}(d)=\gamma_{L}^{*}$ if and only if $A^{L}(d,j_{L}^{*},\gamma_{L}^{*})\hat{p}\leq c^{L}(d,j_{L}^{*},\gamma_{L}^{*})$, where $j_{L}^{*}\in\{1,\ldots,J_{L}\}$ and $\gamma_{L}^{*}$ is in the support of $\hat{\gamma}_{L}(d)$.
@@
    For some finite-valued random vector $\hat{\gamma}_{U}(d)$, some fixed and known matrix $A^{U}(d,j_{U}^{*},\gamma_{U}^{*})$ and some fixed and known vector $c^{U}(d,j_{U}^{*},\gamma_{U}^{*})$, $d\in\widehat{\mathcal{D}}$, $\hat{j}_{U}(d)=j_{U}^{*}$ and $\hat{\gamma}_{U}(d)=\gamma_{U}^{*}$ if and only if $A^{U}(d,j_{U}^{*},\gamma_{U}^{*})\hat{p}\leq c^{U}(d,j_{U}^{*},\gamma_{U}^{*})$, where $j_{U}^{*}\in\{1,\ldots,J_{U}\}$ and $\gamma_{U}^{*}$ is in the support of $\hat{\gamma}_{U}(d)$.
Although not immediately obvious, this assumption holds in a variety of settings; see the examples below. In many cases, this assumption can be simplified because $d\in\widehat{\mathcal{D}}$ if and only if $A_{\mathcal{D}}\hat{p}\leq c_{\mathcal{D}}$ for some fixed and known matrix $A_{\mathcal{D}}$ and vector $c_{\mathcal{D}}$. For these cases, $\hat{\gamma}_{L}(d)$ and $\hat{\gamma}_{U}(d)$ can be vacuously set to fixed constants and
$A^{L}(d,j,\gamma)=\left(\begin{array}[]{c}A_{L}(d,j)\\ A_{\mathcal{D}}\end{array}\right),\quad A^{U}(d,j,\gamma)=\left(\begin{array}[]{c}A_{U}(d,j)\\ A_{\mathcal{D}}\end{array}\right)$
@@
$c^{L}(d,j,\gamma)=\left(\begin{array}[]{c}c_{L}(d,j)\\ c_{\mathcal{D}}\end{array}\right)\quad c^{U}(d,j,\gamma)=\left(\begin{array}[]{c}c_{U}(d,j)\\ c_{\mathcal{D}}\end{array}\right)$
for $A_{L}(d,j)=(\ell_{d,1}-\ell_{d,j},\ldots,\ell_{d,J_{L}}-\ell_{d,j})^{\prime}$, $A_{U}(d,j)=(u_{d,j}-u_{d,1},\ldots,u_{d,j}-u_{d,J_{L}})^{\prime}$, $c_{L}(d,j)=(\tilde{\ell}_{d,j}-\tilde{\ell}_{d,1},\ldots,\tilde{\ell}_{d,j}-\tilde{\ell}_{d,J_{L}})^{\prime}$ and $c_{U}(d,j)=(\tilde{u}_{d,1}-\tilde{u}_{d,j},\ldots,\tilde{u}_{d,J_{L}}-\tilde{u}_{d,j})^{\prime}$. A leading example of this special case is
$\widehat{\mathcal{D}}=\left\{d\in\{0,\ldots,T\}:\widehat{U}(d)\geq\max_{d\in\{0,\ldots,T\}}\widehat{L}(d)\right\},$
@@
for all $d^{\prime}\in\{0,\ldots,T\}$, $j\in\{1,\ldots,J_{U}\}$ and $j^{\prime}\in\{1,\ldots,J_{L}\}$.
We also note that Assumption 3.2 is compatible with the absence of data-dependent selection for which the researcher is interested in forming a CI for an identified interval $[L(d^{*}),U(d^{*})]$ chosen by the researcher a priori. In these cases, $\widehat{\mathcal{D}}=\{d^{*}\}$ and $\hat{\gamma}_{M}(d^{*})$ can be vacuously set to a fixed constant, $A^{M}(d^{*},j,\gamma)=A_{M}(d^{*},j)$ and $c^{M}(d^{*},j,\gamma)=c_{M}(d^{*},j)$ for $M=L,U$. Indeed, we examine an example of this special case when conducting a finite-sample power comparison in Section 7 below.
In general, less conditioning is more desirable in terms of the lengths of the CIs we propose. Although conditioning on the events $d\in\widehat{\mathcal{D}}$, $\hat{j}_{L}(d)=j_{L}^{*}$ and $\hat{j}_{U}(d)=j_{U}^{*}$ is necessary to construct our CIs (see Section 5 below), the researcher should therefore minimize the number of elements in $\hat{\gamma}_{L}(d)$ and $\hat{\gamma}_{U}(d)$ subject to satisfying Assumption 3.2, when constructing our CIs. In some cases it is necessary to condition on these additional random vectors in order to satisfy Assumption 3.2. But in many cases, such as the example given immediately above, additional conditioning random vectors are unnecessary and can be vacuously set to fixed constants.
@@
The notation of this assumption makes explicit that the parameter $p$ and the asymptotic variance $\Sigma$ depend upon the unknown distribution of the data $\mathbb{P}$. It holds naturally for standard estimators $\hat{p}$ under random sampling or weak dependence in the presence of bounds of the moments and dependence of the underlying data.
@@
We impose one final assumption for our unconditional hybrid confidence intervals in order for the object of inferential interest to be well-defined unconditionally.
Assumption 3.6.
$\widehat{\mathcal{D}}=\{\hat{d}\}$ almost surely for a random variable $\hat{d}$ with support $\{0,\ldots,T\}$.
In conjunction, Assumptions 3.2 and 3.6 hold naturally when the object of interest $\hat{d}$ is selected by uniquely maximizing a linear combination of the estimates of the bounds characterizing the identified intervals and the additional conditioning vectors $\hat{\gamma}_{L}(d)$ and $\hat{\gamma}_{U}(d)$ are defined appropriately. Leading examples of this form of selection include when $\hat{d}$ corresponds to the largest estimated lower bound, upper bound or weighted average of lower and upper bounds.
Proposition 3.1.
Suppose $\widehat{\mathcal{D}}=\{\hat{d}\}$, where $\hat{d}=\operatorname*{argmax\;}_{d\in\{0,\ldots,T\}}\{w_{L}\widehat{L}(d)+w_{U}\widehat{U}(d)\}$ is unique almost surely for some fixed known weights $w_{L},w_{U}\geq 0$. Then Assumptions 3.2 and 3.6 are satisfied for
  1. 1.
  2. 2.
  3. 3.
As this proposition makes clear, the additional conditioning vectors needed for Assumption 3.6 to hold depend upon the particular form of selection rule used by the researcher. For example, when $\hat{d}$ is chosen to maximize the estimated lower bound of the identified set $\widehat{L}(d)$, one must condition not only on the realized value of $\hat{j}_{U}(\hat{d})$ when forming a probabilistic upper bound for $U(\hat{d})$ but also $\hat{j}_{L}(\hat{d})$. On the other hand, the formation of either a probabilistic lower bound for $L(\hat{d})$ or upper bound for $U(\hat{d})$ when $\hat{d}$ is chosen to maximize the estimated upper bound of the identified set $\widehat{U}(d)$ requires conditioning on the entire vector $(\hat{j}_{U}(0),\ldots,\hat{j}_{U}(T))^{\prime}$.
Although intuitively appealing, the treatment choice rules of the form described in Proposition 3.1 can be sub-optimal from a statistical decision-theoretic point of view (see, e.g., Manski,, 2021, 2023 and Christensen et al.,, 2023). In Section 4.6, we show how proper definition of $\hat{\gamma}_{L}(d)$ and $\hat{\gamma}_{U}(d)$ satisfies Assumptions 3.2 and 3.6 in the context of optimal selection rules.
@@ 4 Examples
In this section, we show that the proposed inference method is applicable to various examples for which parameters are interval-identified. In particular, we show that Assumptions 3.13.3 are satisfied in these examples.
4.1 Revisiting Manski Bounds with a Continuous Outcome
We revisit the example with Manski bounds in Section 2, now allowing for a continuous outcome with bounded support. Let $Y\in[y^{l},y^{u}]$ be continuously distributed and suppose $E[Y(d)|Z]=E[Y(d)]$ for $d\in\{0,1\}$. Note that
$L(d)$ $\equiv\max_{z\in\{0,1\}}[E[Y|D=d,Z=z]P(D=d|Z=z)+(1-P(D=d|Z=z))y^{l}],$
$U(d)$ $\equiv\min_{z\in\{0,1\}}[E[Y|D=d,Z=z]P(D=d|Z=z)+(1-P(D=d|Z=z))y^{u}],$
and $L(1,0)$ and $U(1,0)$ are Manski’s sharp bounds on the ATE for
$L(\tilde{d},d)$ $\equiv L(\tilde{d})-U(d),$
$U(\tilde{d},d)$ $\equiv U(\tilde{d})-L(d).$
We can define the identified set $\mathcal{D}^{*}\subseteq\mathcal{D}$ of optimal treatments as
$\mathcal{D}^{*}$$\equiv\{d\in\{0,1\}:L(\tilde{d},d)\leq 0,\forall\tilde{d}\in\{0,1\}\}=\{d\in\{0,1\}:\max_{\tilde{d}\in\{0,1\}}L(\tilde{d})\leq U(d)\}.$
Then, Assumption 3.1 holds with
$p$ $=\left(\begin{array}[]{c}E(Y|D=0,Z=1)P(D=0|Z=1)\\ E(Y|D=0,Z=0)P(D=0|Z=0)\\ E(Y|D=1,Z=1)P(D=1|Z=1)\\ E(Y|D=1,Z=0)P(D=1|Z=0)\\ P(D=0|Z=1)\\ P(D=0|Z=0)\end{array}\right)$
and
~$\tilde{\ell}_{0,j}$ $=y^{l}\text{ for }j\in\{0,1\},$
~$\tilde{\ell}_{1,j}$ $=0\text{ for }j\in\{0,1\},$
$\ell_{0,0}$ $=\left(\begin{array}[]{cccccc}0&1&0&0&0&-y^{l}\end{array}\right),$
$\ell_{0,1}$ $=\left(\begin{array}[]{cccccc}1&0&0&0&-y^{l}&0\end{array}\right),$
$\ell_{1,0}$ $=\left(\begin{array}[]{cccccc}0&0&0&1&0&y^{l}\end{array}\right),$
$\ell_{1,1}$ $=\left(\begin{array}[]{cccccc}0&0&1&0&y^{l}&0\end{array}\right)$
$\widehat{\mathcal{D}}$ =$=\{d\in\{0,1\}:\max_{\tilde{d}\in\{0,1\}}\widehat{L}(\tilde{d})\leq\widehat{U}(d)\}.$
For Assumption 3.3, let $[y^{l},y^{u}]=[0,1]$ for simplicity. Note that
[…]

### v2

@@ 3 General Inference Framework
We now introduce the general inference framework that we propose, nesting the Manski bound example of the previous section as a special case. After introducing the general framework, we describe several additional example applications that fall within this framework.
We are interested in performing inference on a parameter $W(d)$ that is indexed by a finite set $d\in\mathcal{D}\equiv\{d^{0},\ldots,d^{K}\}$ for some $K>0$. The index $d$ may correspond to a particular treatment, treatment allocation rule or policy, depending upon the application. We assume that $W(d)$ belongs to an identified set taking a particular interval form that is common to many applications of interest.
@@
For all $d\in\{d^{0},\ldots,d^{K}\}$ and an unknown finite-dimensional parameter $p$,
@@
    $L(d)\equiv\max_{j\in\{1,\ldots,J_{L}\}}\{\tilde{\ell}_{d,j}+\ell_{d,j}p\}\leq W(d)$ for some fixed and known $J_{L}$, ~$\tilde{\ell}_{d,1},\ldots,\tilde{\ell}_{d,J_{L}}$ and nonzero row vectors $\ell_{d,1},\ldots,\ell_{d,J_{L}}$ such that $\ell_{d,j}\neq\ell_{d,j^{\prime}}$ for $j\neq j^{\prime}$.
@@
    $U(d)\equiv\min_{j\in\{1,\ldots,J_{U}\}}\{\tilde{u}_{d,j}+u_{d,j}p\}\geq W(d)$ for some fixed and known $J_{U}$, $\tilde{u}_{d,1},\ldots,\tilde{u}_{d,J_{U}}$ and nonzero row vectors $u_{d,1},\ldots,u_{d,J_{U}}$ such that $u_{d,j}\neq u_{d,j^{\prime}}$ for $j\neq j^{\prime}$.
The lower and upper endpoints of identified sets for the welfare, average potential outcome or ATE typically take the form of $L(d)$ and $U(d)$, especially when (sequences of) outcomes, treatments and instruments are discrete.
In the setting of this paper, a researcher’s interest in $W(d)$ arises when $d$ belongs to a set $\widehat{\mathcal{D}}\subset\mathcal{D}$ that is estimated from a sample of $n$ observations. It is often the case that $\widehat{\mathcal{D}}$ is an estimate of the identified set of best performers $\mathcal{D}^{*}$. This set could correspond to an estimated set of optimal treatments or policies or other data-dependent index sets of interest. The estimated set is determined by an estimator $\hat{p}$ of the finite-dimensional parameter $p$ that determines the bounds on $W(d)$ according to Assumption 3.1. Let
$\hat{j}_{L}(d)\equiv\operatorname*{argmax\;}_{j\in\{1,\ldots,J_{L}\}}\{\tilde{\ell}_{d,j}+\ell_{d,j}\hat{p}\},\quad\hat{j}_{U}(d)\equiv\operatorname{argmin}_{j\in\{1,\ldots,J_{U}\}}\{\tilde{u}_{d,j}+u_{d,j}\hat{p}\},$ (3.1)
which are the indices at which the estimated lower and upper bounds are realized. Then, the estimated lower and upper bounds for $W(d)$ are equal to ~$\tilde{\ell}_{d,\hat{j}_{L}(d)}+\ell_{d,\hat{j}_{L}(d)}\hat{p}$ and $\tilde{u}_{d,\hat{j}_{U}(d)}+u_{d,\hat{j}_{U}(d)}\hat{p}$. We work under the high-level assumption that the following event can be written as a polyhedron in $\hat{p}$: (i) an option index $d$ is in the set of interest $\widehat{\mathcal{D}}$, (ii) the estimated bounds on $W(d)$ are realized at a given value and (iii) (optionally) an additional random vector is realized at any given value.
@@
    For some fixed and known matrix $A^{L}(d,j_{L}^{*},\gamma_{L}^{*})$, some fixed and known vector $c^{L}(d,j_{L}^{*},\gamma_{L}^{*})$ and some finite-valued random vector $\hat{\gamma}_{L}(d)$, the event $\{d\in\widehat{\mathcal{D}}$, $\hat{j}_{L}(d)=j_{L}^{*}$ and $\hat{\gamma}_{L}(d)=\gamma_{L}^{*}\}$ is equivalent to $\{A^{L}(d,j_{L}^{*},\gamma_{L}^{*})\hat{p}\leq c^{L}(d,j_{L}^{*},\gamma_{L}^{*})\}$, where $j_{L}^{*}\in\{1,\ldots,J_{L}\}$ and $\gamma_{L}^{*}$ is in the support of $\hat{\gamma}_{L}(d)$.
@@
    For some fixed and known matrix $A^{U}(d,j_{U}^{*},\gamma_{U}^{*})$, some fixed and known vector $c^{U}(d,j_{U}^{*},\gamma_{U}^{*})$ and some finite-valued random vector $\hat{\gamma}_{U}(d)$, the event $\{d\in\widehat{\mathcal{D}}$, $\hat{j}_{U}(d)=j_{U}^{*}$ and $\hat{\gamma}_{U}(d)=\gamma_{U}^{*}\}$ is equivalent to $\{A^{U}(d,j_{U}^{*},\gamma_{U}^{*})\hat{p}\leq c^{U}(d,j_{U}^{*},\gamma_{U}^{*})\}$, where $j_{U}^{*}\in\{1,\ldots,J_{U}\}$ and $\gamma_{U}^{*}$ is in the support of $\hat{\gamma}_{U}(d)$.
Depending upon the application, $\hat{\gamma}_{L}(d)$ and $\hat{\gamma}_{U}(d)$ (and thus $\gamma_{L}^{*}$ and $\gamma_{U}^{*}$) in this assumption may not be necessary to condition on, in which case they can be vacuously set to constants. Although not immediately obvious, this assumption holds in a variety of settings; see the examples below. In many cases, this assumption can be simplified because, consistent with Assumption 3.1, $d\in\widehat{\mathcal{D}}$ if and only if $A_{\mathcal{D}}\hat{p}\leq c_{\mathcal{D}}$ for some fixed and known matrix $A_{\mathcal{D}}$ and vector $c_{\mathcal{D}}$. For these cases, $\hat{\gamma}_{L}(d)$ and $\hat{\gamma}_{U}(d)$ are not needed and can be vacuously set to fixed constants and
$A^{L}(d,j,\gamma)$ $=\left(\begin{array}[]{c}\ell_{d,1}-\ell_{d,j}\\ \vdots\\ \ell_{d,J_{L}}-\ell_{d,j}\\ A_{\mathcal{D}}\end{array}\right),\quad A^{U}(d,j,\gamma)=\left(\begin{array}[]{c}u_{d,j}-u_{d,1}\\ \vdots\\ u_{d,j}-u_{d,J_{L}}\\ A_{\mathcal{D}}\end{array}\right)$ (3.10)
@@
$c^{L}(d,j,\gamma)$ =(~$=\left(\begin{array}[]{c}\tilde{\ell}_{d,j}-\tilde{\ell}_{d,1}\\ \vdots\\ \tilde{\ell}_{d,j}-\tilde{\ell}_{d,J_{L}}\\ c_{\mathcal{D}}\end{array}\right),\quad c^{U}(d,j,\gamma)=\left(\begin{array}[]{c}\tilde{u}_{d,1}-\tilde{u}_{d,j}\\ \vdots\\ \tilde{u}_{d,J_{L}}-\tilde{u}_{d,j}\\ c_{\mathcal{D}}\end{array}\right).$ (3.19)
A leading example of this special case is
$\widehat{\mathcal{D}}=\left\{d\in\{d^{0},\ldots,d^{K}\}:\widehat{U}(d)\geq\max_{d\in\{d^{0},\ldots,d^{K}\}}\widehat{L}(d)\right\},$
@@
for all $d^{\prime}\in\{d^{0},\ldots,d^{K}\}$, $j\in\{1,\ldots,J_{U}\}$ and $j^{\prime}\in\{1,\ldots,J_{L}\}$.
We also note that Assumption 3.2 is compatible with the absence of data-dependent selection for which the researcher is interested in forming a CI for an identified interval $[L(d^{*}),U(d^{*})]$ chosen by the researcher a priori. In these cases, $\widehat{\mathcal{D}}=\{d^{*}\}$, $\hat{\gamma}_{M}(d^{*})$ can be vacuously set to a fixed constant, $A^{M}(d^{*},j,\gamma)=A_{M}(d^{*},j)$ and $c^{M}(d^{*},j,\gamma)=c_{M}(d^{*},j)$ for $M=L,U$. Indeed, we examine an example of this special case when conducting a finite-sample power comparison in Section 6 below.
In general, less conditioning is more desirable in terms of the lengths of the CIs we propose. Although conditioning on the events $d\in\widehat{\mathcal{D}}$, $\hat{j}_{L}(d)=j_{L}^{*}$ and $\hat{j}_{U}(d)=j_{U}^{*}$ is necessary to construct our CIs (see Section 5.1 below), the researcher should therefore minimize the number of elements in $\hat{\gamma}_{L}(d)$ and $\hat{\gamma}_{U}(d)$ subject to satisfying Assumption 3.2 when constructing our CIs. In some cases it is necessary to condition on these additional random vectors in order to satisfy Assumption 3.2. But in many cases, such as the example given immediately above, additional conditioning random vectors are unnecessary and can be vacuously set to fixed constants.
We impose the following assumption for our unconditional hybrid CIs in order for the object of inferential interest to be well-defined unconditionally.
Assumption 3.3.
$\widehat{\mathcal{D}}=\{\hat{d}\}$ almost surely for a random variable $\hat{d}$ with support $\{d^{0},\ldots,d^{K}\}$.
In conjunction, Assumptions 3.2 and 3.3 hold naturally when the object of interest $\hat{d}$ is selected by uniquely maximizing a linear combination of the estimates of the bounds characterizing the identified intervals and the additional conditioning vectors $\hat{\gamma}_{L}(d)$ and $\hat{\gamma}_{U}(d)$ are defined appropriately. Leading examples of this form of selection include when $\hat{d}$ corresponds to the largest estimated lower bound, upper bound or weighted average of lower and upper bounds.
Proposition 3.1.
Suppose $\widehat{\mathcal{D}}=\{\hat{d}\}$, where $\hat{d}=\operatorname*{argmax\;}_{d\in\{d^{0},\ldots,d^{K}\}}\{w_{L}\widehat{L}(d)+w_{U}\widehat{U}(d)\}$ is unique almost surely for some fixed known weights $w_{L},w_{U}\geq 0$. Then Assumptions 3.2 and 3.3 are satisfied for
  1. 1.
  2. 2.
  3. 3.
Expressions for $A^{M}(d,j_{M}^{*},\gamma_{M}^{*})$ and $c^{M}(d,j_{M}^{*},\gamma_{M}^{*})$ for $M=L,U$ in the settings of Proposition 3.1 are available for reference in its proof in Appendix C. As this proposition makes clear, the additional conditioning vectors needed for Assumption 3.3 to hold depend upon the particular form of selection rule used by the researcher. For example, when $\hat{d}$ is chosen to maximize the estimated lower bound of the identified set $\widehat{L}(d)$, one must condition not only on the realized value of $\hat{j}_{U}(\hat{d})$ when forming a probabilistic upper bound for $U(\hat{d})$ but also $\hat{j}_{L}(\hat{d})$. On the other hand, the formation of either a probabilistic lower bound for $L(\hat{d})$ or upper bound for $U(\hat{d})$ when $\hat{d}$ is chosen to maximize the estimated upper bound of the identified set $\widehat{U}(d)$ requires conditioning on the entire vector $(\hat{j}_{U}(0),\ldots,\hat{j}_{U}(T))^{\prime}$.
Although intuitively appealing, the treatment choice rules of the form described in Proposition 3.1 can be sub-optimal from a statistical decision-theoretic point of view (see, e.g., Manski,, 2021, 2023 and Christensen et al.,, 2023). In Section 4.4, we show how proper definition of $\hat{\gamma}_{L}(d)$ and $\hat{\gamma}_{U}(d)$ satisfies Assumptions 3.2 and 3.3 in the context of the optimal selection rules of Christensen et al., (2023).
@@
The notation of this assumption makes explicit that the parameter $p$ and the asymptotic variance $\Sigma$ depend upon the unknown distribution of the data $\mathbb{P}$. It holds naturally for standard estimators $\hat{p}$ under random sampling or weak dependence in the presence of bounds on the moments and dependence of the underlying data.
@@ 4 Examples
In this section, we show that the proposed inference method is applicable to various examples for which parameters are interval-identified. In particular, we show that Assumptions 3.1, 3.2 and 3.4 are satisfied in these examples. See Appendix A for additional examples.